# On robust overfitting: adversarial training induced distribution matters

### 28 Nov 2023 v1

@@ 5 Generalization properties of the induced distributions
For simplicity we rewrite this quantity as $\tilde{\gamma}_{\theta}(x,y)$ and refer to it as the local dispersion of the perturbation (family) ${\cal Q}_{\theta}$ at $(x,y)$.
Notably, for any given $(x,y)$, it is possible to estimate this quantity by Monte-Carlo approximation. In our experiments, to estimate local dispersion $\tilde{\gamma}_{\theta}(x,y)$, we sample 250 pairs of $(\rho,\rho^{\prime})$ and approximate the expectation by the sample mean.
We now show that the expected value of local dispersion of ${\cal Q}_{\theta}$ is a governing term of the generalization gap, under the following conditions.
@@
  • (Boundedness of perturbation-smoothed loss) $\sup\limits_{x,y\in{\rm supp}(\mathcal{D})}\left|\mathbb{E}_{\rho}f(\mathcal{Q}_{x,y,\theta}(x+\rho),y)\right|=A<\infty$, where ${\rm supp}(\mathcal{D})$ denote the support of distribuion ${\cal D}$.
@@
Theorem 1. Let $f\in{\cal F}$ and suppose that $f$ satisfies the above conditions. Then for any $\tau>0$, with probability at least $1-\tau$ over the i.i.d. draws of sample $\{(v_{i},y_{i})\}_{i=1}^{m}$ from $\tilde{\mathcal{D}}_{\theta}$,
@@
$|\frac{1}{m}\sum\limits_{i=1}^{m}f(v_{i},y_{i})-\mathbb{E}f(v,y)|\leq\frac{2\beta}{\sqrt{m}}\sqrt{\mathbb{E}_{{\mathcal{D}^{*}}}\tilde{\gamma}_{\theta}(x,y)}+\frac{2A}{\sqrt{m}}+2B\sqrt{\frac{\log\frac{1}{\tau}}{2m}}$ (8)
@@
We leave the proof in Appendix B. The theorem shows that the generalization gap of any $f$ w.r.t to the distribution $\tilde{\mathcal{D}}_{\theta}$ is affected by the expected local dispersion (ELD) $\mathbb{E}_{{\mathcal{D}^{*}}}\tilde{\gamma}_{\theta}(x,y)$ of ${\cal Q}_{\theta}$. More specifically, it suggests that a small generalization gap can be achieved when $\mathbb{E}_{{\mathcal{D}^{*}}}\tilde{\gamma}_{\theta}(x,y)$ is small.
Figure 4: Experiments on the CIFAR-10 testing set: (a) histograms of $\tilde{\gamma}_{t}(x,y)$ at three distinct PGD-AT checkpoints. (b) The evolution of ELD w.r.t $t$ and the IDE testing error for each $\tilde{\mathcal{D}}_{t}$. The results show that the level of $\tilde{\gamma}_{t}(x,y)$ increases during PGD-AT and correspondingly the model obtained from $\tilde{\mathcal{D}}_{t}$ becomes harder to generalize. This implies that the local property of $\mathcal{Q}_{x,y,\theta_{t}}$ characterized by $\tilde{\gamma}_{t}(x,y)$ plays a dominate role in influencing the generalization of $\tilde{\mathcal{D}}_{t}$.
@@
To validate the usefulness of this theorem, we performed experiments to estimate the ELD for $\theta$ at various checkpoints $t$. Note that the expectation here is over the distribution ${\cal D}^{*}$, from which no samples are available. However, due to the relationship between ${\cal D}_{\cal X}$ and ${\cal D}^{*}_{\cal X}$, namely that ${\cal D}_{\cal X}$ is merely an $\epsilon$-smoothed version of ${\cal D}^{*}_{\cal X}$, one expects that when we draw $x$ from ${\cal D}$, ${\cal D}_{\cal X}(x)\approx{\cal D}^{*}_{\cal X}(x)$ with high probability. Then when we estimate ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{\theta}(x,y)$ using an i.i.d. sample $S$ drawn from ${\cal D}$, we may approximately regard $S$ as drawn i.i.d. from ${\cal D}^{*}$, and estimate ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{\theta}(x,y)$ by ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{\theta}(x,y)\approx\frac{1}{m}\sum_{i=1}^{m}\tilde{\gamma}_{\theta}(x_{i},y_{i}),$ where $\{(x_{i},y_{i})\}_{i=1}^{m}$ are drawn from ${\cal D}$. Notably this is a slightly biased estimator of ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{\theta}(x,y)$, and the bias decreases with $\epsilon$.
@@
In our experiments, we inspect the evolution of the distribution of $\tilde{\gamma}_{\theta_{t}}(x,y)$ (or $\tilde{\gamma}_{t}(x,y)$ for simplicity) along the PGD-AT trajectory. Figure 4(a) plots the histogram of $\tilde{\gamma}_{t}(x,y)$ for the testing set of CIFAR-10 at three different PGD-AT checkpoints. It is clear that the distribution shifts to the right as PGD-AT proceeds, indicating that the perturbation operator ${\cal Q}_{\theta_{t}}$ becomes more locally dispersive as training goes on.
@@
We then estimate ELD on the testing set (the green curve in Figure 4(b)). We plot one of the IDE results using red curve in the same figure for a clearer comparison. As shown in the figure, ELD is getting larger along PGD-AT, correlating with the increasing difficulty of generalization on $\tilde{\mathcal{D}}_{t}$. These results suggests that the bound in Theorem 1 adequately characterizes the generalization behaviour in the IDE experiments, which in turn suggests that the local dispersiveness of the perturbation operator contributes to robust overfitting in PGD-AT. Similar experimental results are also observed across different datasets (see Appendix C).
6 Other implications
@@
Our preceding theoretical analysis underscores the critical role played by the local properties of $\mathcal{Q}_{x,y,\theta_{t}}$ in affecting the generalization performance for $\tilde{\mathcal{D}}_{t}$. This, in turn, inspires our curiosity to investigate whether additional local properties, beyond local dispersion, also posses critical influences on the generalization of $\tilde{\mathcal{D}}_{t}$. As such, we inspect the expected distance between the adversarial examples generated by PGD and its clean counterparts, defined as
$d_{\theta}(x,y):=\mathbb{E}_{\rho\sim\mathcal{U}([-\epsilon,\epsilon]^{d})}\|\mathcal{Q}_{x,y,\theta}(x+\rho)-x\|_{2}$ (9)
By triangle inequality, we notice that
$\mathbb{E}_{\rho,\rho^{\prime}\sim\mathcal{U}([-\epsilon,+\epsilon]^{d})}\|\mathcal{Q}_{x,y,\theta}(x+\rho)-\mathcal{Q}_{x,y,\theta}(x+\rho^{\prime})\|_{2}$ (10)
$=$ $\mathbb{E}_{\rho,\rho^{\prime}\sim\mathcal{U}([-\epsilon,+\epsilon]^{d})}\|\mathcal{Q}_{x,y,\theta}(x+\rho)-x+x-\mathcal{Q}_{x,y,\theta}(x+\rho^{\prime})\|_{2}$ (11)
$\leq$ $2d_{\theta}(x,y)$ (12)
The term (10) is related to the local dispersion of $\mathcal{Q}_{x,y,\theta}$ despite that the definition of the local dispersion computes the square of the $l_{2}$-distance. Recall that we have observed an increase in the level of local dispersion along the PGD-AT trajectory. According to the inequality (12), one might logically expect that the level of $d_{\theta}(x,y)$ should also increase, meaning that the perturbed data generated by $x$ are getting not only more “dispersed” around $x$ but also move farther from $x$. However, our experimental findings present a contradictory result. Instead of an increase, we observe a decrease in the level of $d_{\theta}(x,y)$ during PGD-AT. This unexpected trend suggests that the perturbed data generated by $x$ are, in fact, moving closer to the original data point $x$.
@@
In our experiments, we estimate $d_{\theta}(x,y)$ by computing the sample mean with 250 samples of $\rho$ drawn from $\mathcal{U}([-\epsilon,\epsilon]^{d})$. We analyze the dynamic behavior of $d_{\theta_{t}}(x,y)$, which we refer to as $d_{t}(x,y)$ for simplicity, along the PGD-AT trajectory. In Figure 5 (a), we present histogram of $d_{t}(x,y)$ for the CIFAR-10 testing set at three distinct training checkpoints. Notably, the histogram exhibits a notable mode shift towards a smaller value, indicating a trend that as PGD-AT proceeds, the generated adversarial examples progressively approach their clean counterparts. The reduction in the level of $d_{t}(x,y)$ along PGD-AT is further observed by evaluating the expectation $\mathbb{E}_{\mathcal{D}}d_{t}(x,y)$ on the testing set (see Figure 5 (c), green curve), where a clear drop in $\mathbb{E}_{\mathcal{D}}d_{t}(x,y)$ is exhibited.
As a reminder, we previously noted that adversarial examples tend to become more dispersed around their clean counterparts as training progresses. The experimental findings presented here shed light on this phenomenon, suggesting that the growing dispersion is likely to be a result of the perturbation angles expanding, while the perturbation magnitudes seem to have a lesser impact on the level of dispersion.
Figure 5: Experiments on the CIFAR-10 testing set. (a) and (b): histograms of $d_{t}(x,y)$ and $\Phi_{t}(x,y)$ at different PGD-AT epochs $t$. (c): The evolutions of $\mathbb{E}_{\mathcal{D}}d_{t}(x,y)$ and $\mathbb{E}_{\mathcal{D}}\Phi_{t}(x,y)$ along PGD-AT trajectory. Combined with the results in Figure 4, an interesting phenomenon in PGD-AT is revealed: as the training proceeds, the perturbed data generated by $x$ are getting closer to $x$ and in the meanwhile getting more dispersed potentially due to the spreading of perturbation angles.
To verify this conjecture, we evaluate the expected angle between a pair of perturbations generated from $(x,y)$, defined as
@@
$\Phi_{\theta}(x,y):=\mathbb{E}_{\rho,\rho^{\prime}\sim\mathcal{U}([-\epsilon,+\epsilon]^{d})}\cos^{-1}\left(\frac{(\mathcal{Q}_{x,y,\theta}(x+\rho)-x)^{T}(\mathcal{Q}_{x,y,\theta}(x+\rho^{\prime})-x)}{\|\mathcal{Q}_{x,y,\theta}(x+\rho)-x\|_{2}\|\mathcal{Q}_{x,y,\theta}(x+\rho^{\prime})-x\|_{2}}\right)$ (13)
with the expectation estimated by computing the sample mean of 250 pairs of $\rho,\rho^{\prime}$ drawn from $\mathcal{U}([-\epsilon,+\epsilon]^{d})$. Figure 5 (b) plots the histograms of $\Phi_{\theta_{t}}(x,y)$ (or $\Phi_{t}(x,y)$) at three distinct checkpoint $t$ and Figure 5 (c) illustrate the evolution of $\mathbb{E}_{\mathcal{D}}\Phi_{t}(x,y)$ with the yellow curve. The results present an increase in the level of $\Phi_{t}(x,y)$, indicating a spreading of perturbation angles and less “aligned” perturbations generated by each $x$ during PGD-AT. Similar experimental results have been observed across other datasets. (see Appendix D)
@@
We conjecture that this wider spread of angles in adversarial perturbations is a consequence of an intricate or ”ragged” shape in the model’s decision boundary. In essence, the shape of the decision boundary has a substantial influence on the direction of perturbations. For instance, the perturbations generated by linear classifiers are always aligned due to that the decision boundary is ”smooth”. Conversely, one would expect that a jagged or irregular decision boundary could result in perturbations that are both more dispersed and less aligned. We speculate that the presented dynamics of $\mathbb{E}_{\mathcal{D}}\Phi_{t}(x,y)$ and $\mathbb{E}_{\mathcal{D}}\tilde{\gamma}_{t}(x,y)$ can be explained as: during the early stages of PGD-AT, the adversarial perturbations generated by the data $(x,y)$ exhibit a higher degree of alignment due to the initial smoothness of the model’s decision boundary. This results in a smaller level of $\mathbb{E}_{\mathcal{D}}\Phi_{t}(x,y)$ and $\mathbb{E}_{\mathcal{D}}\tilde{\gamma}_{t}(x,y)$. However, as training progresses, the decision boundary is twisted, in order to fit or ”memorize” the training data, causing the perturbations to become less aligned and more dispersed, leading to the rise in the level of $\mathbb{E}_{\mathcal{D}}\Phi_{t}(x,y)$ and $\mathbb{E}_{\mathcal{D}}\tilde{\gamma}_{t}(x,y)$. Consequently, the increasing dispersion or spread of angles may serve as an indicative measure for the degree of irregularity present in the decision boundary.
@@
Our observation of the increasing dispersion and spreading angles of adversarial perturbations along PGD-AT is, to our best knowledge, a novel finding. This discovery may provide valuable insights into comprehending the dynamic of PGD-AT and the phenomenon of robust overfitting.

### 10 Feb 2024 v2

@@ 5 Generalization properties of the induced distributions
We refer to this quantity as the local dispersion of ${\cal Q}_{x,y,\theta_{t}}$, as it measures the spread of the perturbation of a random point in ${\mathbb{B}}(x,\epsilon)$ by ${\cal Q}_{x,y,\theta_{t}}$ —To see this, one may verify that $\tilde{\gamma}_{t}(x,y)$ can be expressed as
$\tilde{\gamma}_{t}(x,y)=2\cdot{\rm Trace}\left({\rm COV}({\cal Q}_{x,y,\theta_{t}}(V))\right)$
where $V$ is drawn from ${\mathbb{U}}(x,\epsilon)$ and ${\rm COV}({\cal Q}_{x,y,\theta_{t}}(V))$ denotes the covariance matrix of the random vector ${\cal Q}_{x,y,\theta_{t}}(V)$.
One may argue intuitively that smaller local dispersion of ${\cal Q}_{\theta_{t}}$ may allow the resulting distribution $\tilde{\cal D}_{t}$ to generalize better: consider an instance $(X,Y)$ drawn from the absolutely robust distribution ${\cal D}^{*}$, and two data points $(X+\rho,Y)$ and $(X+\rho^{\prime},Y)$ (with $\rho$ and $\rho^{\prime}$ drawn independently from ${\mathbb{U}}(0,\epsilon)$. Suppose that $(X+\rho,Y)$ is included in the training set and $(X+\rho^{\prime},Y)$ is included in the testing set. When the local dispersion $\tilde{\gamma}_{t}(X,Y)$ is small, the perturbed version of training point $(X+\rho,Y)$ and that of the testing point $(X+\rho^{\prime},Y)$ (both of which are realizations from $\tilde{\cal D}_{t}$) are close, allowing the latter to behave similarly as the former.
We now rigorously formalize this intuition, under the following assumptions.
@@
Theorem 5.1.
@@
Let $f\in{\cal F}$ and suppose that $f$ satisfies the above conditions. Then for any $\tau>0$, with probability at least $1-\tau$ over the i.i.d. draws of sample $\{(v_{i},y_{i})\}_{i=1}^{m}$ from $\tilde{\mathcal{D}}_{t}$,
@@
${\rm GG}_{t}(f)\leq\frac{2\beta}{\sqrt{m}}\sqrt{\mathbb{E}_{{\mathcal{D}^{*}}}\tilde{\gamma}_{t}(x,y)}+\frac{2\beta\sqrt{d}\epsilon}{\sqrt{m}}+\frac{2B}{\sqrt{m}}\left(\sqrt{\frac{\log\frac{1}{\tau}}{2}}+1\right)$ (8)
@@
The derivation of the bound is based on a modification of the Rademacher complexity analysis. To directly analyze the Rademacher complexity of the hypothesis class ${\cal F}$ is however difficult in our setting due to the lack of restrictions on ${\cal F}$. A future direction towards our work is extending our analysis to more restricted hypothesis class. We leave the proof of the Theorem in Appendix B. The theorem shows that the generalization gap of any $f$ w.r.t to the distribution $\tilde{\mathcal{D}}_{t}$ is affected by the expected local dispersion (ELD) $\mathbb{E}_{{\mathcal{D}^{*}}}\tilde{\gamma}_{t}(x,y)$ of ${\cal Q}_{\theta_{t}}$. The key message of this theorem is that a small generalization gap is guaranteed when $\mathbb{E}_{{\mathcal{D}^{*}}}\tilde{\gamma}_{t}(x,y)$ is small.
@@
To validate this theorem, we performed experiments to estimate the ELD for $\theta_{t}$ at various checkpoints $t$. Note that the expectation here is over the distribution ${\cal D}^{*}$, from which no samples are available. However, due to the relationship between ${\cal D}_{\cal X}$ and ${\cal D}^{*}_{\cal X}$, namely that ${\cal D}_{\cal X}$ is merely an $\epsilon$-smoothed version of ${\cal D}^{*}_{\cal X}$, one expects that when we draw $x$ from ${\cal D}$, ${\cal D}_{\cal X}(x)\approx{\cal D}^{*}_{\cal X}(x)$ with high probability. Then when we estimate ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{t}(x,y)$ using an i.i.d. sample $S$ drawn from ${\cal D}$, we may approximately regard $S$ as drawn i.i.d. from ${\cal D}^{*}$, and estimate ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{t}(x,y)$ by ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{t}(x,y)\approx\frac{1}{m}\sum_{i=1}^{m}\tilde{\gamma}_{t}(x_{i},y_{i}),$ where $\{(x_{i},y_{i})\}_{i=1}^{m}$ are drawn from ${\cal D}$. Notably this is a slightly biased estimator of ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{t}(x,y)$, and the bias decreases with $\epsilon$. In our experiments, the estimation of each $\tilde{\gamma}_{t}(x_{i},y_{i})$ is done by Monte-Carlo approximation via sampling 250 pairs of $(\rho,\rho^{\prime})$ and approximating the expectation in (7) by the sample mean.
@@
We inspect the evolution of the distribution of $\tilde{\gamma}_{t}(x,y)$ along the PGD-AT trajectory. Figure LABEL:fig:_disper (a) and (b) respectively plot the histograms of $\tilde{\gamma}_{t}(x,y)$ for the testing set of CIFAR-10 and CIFAR-100 at three different PGD-AT checkpoints. It is clear that the distribution shifts to the right as PGD-AT proceeds, indicating that the perturbation operator ${\cal Q}_{\theta_{t}}$ becomes more locally dispersive as training goes on.
@@
We then estimate ELD on the testing sets (the green curves in Figure LABEL:fig:_disper (c) and (d)). We plot the corresponding IDE results using red curve in each sub-figure as a clearer comparison. As shown in the figures, ELD is getting larger along PGD-AT, correlating with the increasing difficulty of generalization on $\tilde{\mathcal{D}}_{t}$. These results confirm that the bound in Theorem 8 adequately characterizes the generalization behaviour in the IDE experiments, which in turn suggests that the local dispersiveness of the perturbation operator contributes to robust overfitting in PGD-AT. Similar experimental results are also observed across other datasets (see Appendix C).
6 Other observations and discussions
@@
Our preceding theoretical analysis underscores the critical role played by the local properties of $\mathcal{Q}_{x,y,\theta_{t}}$ in affecting the generalization performance of models learned on samples from $\tilde{\mathcal{D}}_{t}$. This, in turn, inspires our curiosity to investigate whether additional local properties, beyond local dispersion, also posses critical influences on the generalization of $\tilde{\mathcal{D}}_{t}$. As such, we inspect the expected distance between the adversarial examples generated by PGD and its clean counterparts, defined as
$d_{t}(x,y):=\mathbb{E}_{\rho\sim{\mathbb{U}}(0,\epsilon)}\|\mathcal{Q}_{x,y,\theta_{t}}(x+\rho)-x\|_{2}$ (9)
By triangle inequality, it can be easily deduced that when $\rho$ and $\rho^{\prime}$ are drawn independently from ${\mathbb{U}}(0,\epsilon)$,
$\mathbb{E}_{\rho,\rho^{\prime}}\|\mathcal{Q}_{x,y,\theta_{t}}(x+\rho)-\mathcal{Q}_{x,y,\theta_{t}}(x+\rho^{\prime})\|_{2}\leq 2d_{t}(x,y)$ (10)
It is apparent that the term on the left hand side of this inequality is related to the local dispersion of $\mathcal{Q}_{x,y,\theta_{t}}$ despite that the local dispersion is defined using the square of the $l_{2}$-distance. This inequality then motivates us to investigate if $d_{t}(x,y)$ may behave similarly as the local dispersion $\tilde{\gamma}_{t}(x,y)$.
@@
We conducted Monte-Carlo estimation of $d_{t}(x,y)$ by sampling 250 realizations of $\rho$ drawn from ${\mathbb{U}}(0,\epsilon)$. We analyze the dynamic behavior of $d_{t}(x,y)$ along the PGD-AT trajectory. In Figure LABEL:fig:_d2c_angle (a) and (b), we present histogram of $d_{t}(x,y)$ for the CIFAR-10 and CIFAR-100 testing set at three distinct training checkpoints. Notably, the histogram exhibits a notable mode shift towards a smaller value, indicating a trend that as PGD-AT proceeds, the generated adversarial examples progressively approach their clean counterparts. The reduction in the level of $d_{t}(x,y)$ along PGD-AT is further observed by estimating the expectation $\mathbb{E}_{\mathcal{D}^{*}}d_{t}(x,y)$ (again, the expectation over ${\cal D}^{*}$ is approximated by expectation over ${\cal D}$) on the testing set (see Figure LABEL:fig:_d2c_angle (e) and (f), green curves), where a clear drop in $\mathbb{E}_{\mathcal{D}^{*}}d_{t}(x,y)$ is exhibited. From these results, the behavior of $d_{t}(x,y)$ along PGD-AT shows a drastically different trend. Notably, although these results might seem to suggest that there is strong negative correlation between $\gamma_{t}(x,y)$ and $d_{t}(x,y)$, additional observations (see Appendix D Figure LABEL:fig:_disper_d2c) suggest that this need not be the case. Thus ${\mathbb{E}}_{{\cal D}^{*}}d_{t}(x,y)$ is not as indicative of the IDE generalization gap (and hence of robust overfitting) as ${\mathbb{E}}_{{\cal D}^{*}}\tilde{\gamma}_{t}(x,y)$.
To develop a deeper understanding of the difference between $\tilde{\gamma}_{t}(x,y)$ and $\tilde{d}_{t}(x,y)$ and the above observations, consider random instance $(x,y)$ drawn from ${\cal D}^{*}$, which we will refer to as a “robust anchor” —recalling that ${\cal D}^{*}$ is absolutely robust. Draw $\rho$ from ${\mathbb{U}}(0,\epsilon)$, and define random variable $\Delta_{t}(x,y):=\mathcal{Q}_{x,y,\theta_{t}}(x+\rho)-x$. It is clear that $d_{t}(x,y)$ is the mean norm of $\Delta_{t}(x,y)$ and $\tilde{\gamma}_{t}(x,y)$ is (up to scale) the trace of the covariance of $\Delta_{t}(x,y)$, as we shown before. Then we see that $d_{t}(x,y)$ measures, on average, how far the data point $x+\rho$ is away from its robust anchor $x$ after it is perturbed by ${\cal Q}_{x,y,\theta_{t}}$; thus $d_{t}(x,y)$ may be seen as the “bias” that the perturbation operator introduces to a data point relative to its robust anchor. On the other hand, $\gamma_{t}(x,y)$ is a notion of “variance”, measuring, on average, to what extent the data point $x$ is dispersed by the perturbation operator with respect to its robust anchor. The above experimental observations then simply suggest that it is the variance, not the bias, that is closely related to the IDE genearalization gap and to robust overfitting.
As previously noted, when robust overfitting occurs, the local dispersion tends to increase along the PGD-AT trajectory. It is intuitively sensible that increased dispersion may lead to wider spread of the perturbation angles. To formalize this notion of angular spread, we define
$\Phi_{t}(x,y):=\mathbb{E}_{\rho,\rho^{\prime}}\cos^{-1}\left(\frac{Z^{T}_{t}(x,y)Z^{\prime}_{t}(x,y)}{\|Z_{t}(x,y)\|_{2}\|Z_{t}(x,y)\|_{2}}\right)$ (11)
@@
where $\rho$ and $\rho^{\prime}$ are drawn independently from ${\mathbb{U}}(0,\epsilon)$, $Z_{t}(x,y):=\mathcal{Q}_{x,y,\theta_{t}}(x+\rho)$, and $Z^{\prime}_{t}(x,y):=\mathcal{Q}_{x,y,\theta_{t}}(x+\rho^{\prime})$.
Via a similar Monte-Carlo approach, we may estimate $\Phi_{t}(x,y)$ for any point $(x,y)$. Figure LABEL:fig:_d2c_angle (c) and (d) respectively plot the histograms of $\Phi_{t}(x,y)$ on CIFAR-10 and CIFAR-100 at three distinct checkpoints $t$ and the the yellow curves in Figure LABEL:fig:_d2c_angle (e) and (f) illustrate the evolution of $\mathbb{E}_{\mathcal{D}^{*}}\Phi_{t}(x,y)$. The results present an increasing trend of $\Phi_{t}(x,y)$, indicating a widening spread of the perturbation angles as PGD-AT proceed. This observation correlates well with the increasing trend of local dispersion. Similar experimental results have been observed across other datasets. (see Appendix D).
We conjecture that this wider spread of angles is due to the increasingly complex shape of the model’s decision boundary along adversarial training. Notably the shape of the decision boundary has a substantial influence on the direction of adversarial perturbations. For instance, adversarial perturbations against a linear classifiers are always aligned, namely, all perpendicular to the decision hyperplane. As the decision boundary bends to increasingly sharper angles and to more complex shapes, adversarial perturbation on points in a small neigbourhood of a robust anchor can vary greatly in directions.
@@
We speculate that during the early stage of PGD-AT, the model’s decision boundary quickly arrive at a smooth shape, causing adversarial perturbations to exhibit a higher degree of alignment. This results in a smaller level of $\mathbb{E}_{\mathcal{D}^{*}}\Phi_{t}(x,y)$ and $\mathbb{E}_{\mathcal{D}^{*}}\tilde{\gamma}_{t}(x,y)$. However, as training progresses, the decision boundary evolves to a more complex shape, in order to fit or ”memorize” the adversarially perturbed training data. This causing the subsequent adversarial perturbations to become less aligned and more dispersed, resulting in the rise of $\mathbb{E}_{\mathcal{D}^{*}}\Phi_{t}(x,y)$ and $\mathbb{E}_{\mathcal{D}^{*}}\tilde{\gamma}_{t}(x,y)$. Consequently, the increasing dispersion or spread of angles may serve as an indicative measure for the degree of irregularity, or the “shape complexity”, of the decision boundary.
@@
Our observation of the increasing dispersion and spreading angles of adversarial perturbations along PGD-AT is, to our best knowledge, a novel finding. This discovery may provide valuable insights into comprehending the dynamics of PGD-AT and the phenomenon of robust overfitting.